2 moles of mixture of co and co2 requires exactly 56 g of koh in soltion for complete conversion of all the co2 intok2co3.how much amount more of koh will be required for conversion into k2co3.if two moles of mixture is completely oxidised to c02

Grade:

1 Answers

Sunil Kumar FP
9 years ago
WE have
2KOH +CO2 ---------K2CO3 +H2O
2 mole of KOH is requireb one mole of CO2 fromthe above reaction
therefore weight of CO2 required by 56 gm of KOH=1/2 MOLE of CO2=1/2*44=22gm
Since it requires half mole exactly therfore 1.5 is the amount od CO.
CO +1/2O2---CO2
1.5 mole of CO2 will be formed.
thus 2*1.5 mole=3 mole of KOH will be required.more KOH required=3*56 gm=168gm